Please help me to find the first five terms of the following recursively defined infinite sequence.
a1=2 (1 represents the input value of term number)
ak+1=3ak-5 (k+1,k represents the input value of term number)

TQ.

Answer by stanbon(75887)

find the first five terms of the following recursively defined infinite sequence.
a1=2 (1 represents the input value of term number)
ak+1=3ak-5 (k+1,k represents the input value of term number)
----
1st: 2
2nd: 3*2-5 = 1
3rd: 3*1-5 = -2
4th: 3*-2-5 = -11
5th: 3*-11-5 = -38
========================
Cheers,
Stan H.
=============